OpenStudy (anonymous):

Hi everyone! I'm doing an inverse Laplace transform s/(s^2 + 4s + 5)...then I use completing the square to get to s/ [(s+2)^2 + 1]...now this problem asks you to use the first translation theorem L{e^at * f(t)} = F(s-a) so the next step is to add and subtract 2 on the numerator like this >>> (s + 2 - 2)/ [(s+2)^2 + 1]...my question is why do they use 2? Thanks! :o)

OpenStudy (anonymous):

why not +9-9 for example?

OpenStudy (anonymous):

@rational

OpenStudy (anonymous):

@ParthKohli can you help me? :o)

Parth (parthkohli):

Ah, I'm not too familiar with DEs yet, but your question seems to be mostly about algebraic manipulation. I don't know how the theorem is applied, but I can tell that they did \((s+2) -2\) keeping in mind the \((s+2)^2 + 1\) in the denominator.

OpenStudy (anonymous):

i'm trying to figure out how to word my question..one sec lol

OpenStudy (anonymous):

when they did +2-2, it looks like that order is somehow significant when they eventualy split up the problem because the s+2 on top and bottom get separated from the -2 like this>>> [ (s+2)/(s+2)^2+1 ] - [ 2/ )s+2)^2 +1 ]...see what I mean? I wonder if they would have gotten the wrong answer if they put the s-2 in the numerator on the left side ?

Parth (parthkohli):

Yeah, I believe you just answered your own question! Expressions of the form \(\frac{t}{t^2 + 1}\) and \(\frac{1}{t^2 + 1}\) are simple to work with, at least, in calc I. (Here, \(t := s+2\).) I do not know the Laplace Transform part of it, but what I can tell from context is that you wouldn't have been able to do the same with any other kind of splitting.

Parth (parthkohli):

@rational knows this stuff, and he may be able to add to my argument.

OpenStudy (anonymous):

I tagged him but he must be busy

Parth (parthkohli):

@Callisto?

OpenStudy (anonymous):

test today and I need to figure this out

Parth (parthkohli):

Do you know the rest of the solution? Can you see in the later steps how this manipulation was used?

OpenStudy (anonymous):

i wonder...if they had instead (s-9) in the denominator, would they have chosen -9+9 on top and both (s-9) would have traveled off together to the left hand side?

Parth (parthkohli):

Yes, that's correct!

OpenStudy (anonymous):

yes I see what they did, but don't know why

Parth (parthkohli):

If it led to the solution, you do.

OpenStudy (anonymous):

I guess you need to keep the like quantities together because ultimately, that s+2 is actually s-(-2) and a=-2 then

OpenStudy (anonymous):

you need the actual value of "a" in order to get the final solution

OpenStudy (anonymous):

I'm pretty sure that is the reasoning, at least I can't think of anything else

OpenStudy (anonymous):

I will leave this question open for a bit in case someone can confirm what we have written...thanks Parth! :o)

OpenStudy (anonymous):

\[\mathcal{L}\{\sin t\}=\frac{s}{s^2+1};\quad\mathcal{L}\{\cos t\}=\frac1{s^2+1}\] so \[\mathcal{L}\{e^{-at}\sin t\}=\frac{s+a}{(s+a)^2+1};\quad\mathcal{L}\{e^{-at}\cos t\}=\frac1{(s+a)^2+1}\] in our case, we have: $$\begin{align*}\mathcal{L}^{-1}\left\{\frac{s}{(s+2)^2+1}\right\}&=\mathcal{L}^{-1}\left\{\frac{(s+2)-2}{(s+2)^2+1}\right\}\\&=\mathcal{L}^{-1}\left\{\frac{s+2}{(s+2)^2+1}\right\}-2\cdot \mathcal{L}^{-1}\left\{\frac1{(s+2)^2+1}\right\}\\&=e^{-2t}\sin t-2e^{-2t}\cos t\end{align*}$$

OpenStudy (anonymous):

so you MUST keep the (s+2) with the other (s+2) or it will not work correct?

OpenStudy (anonymous):

we split the numerator \(s=(s+2)-2\) in such a manner because we notice the similarity of our expression to that of the Laplace transform of \(e^{-at}\sin t\) with \(a=2\). the purpose of splitting $(s+2)-2$ rather than $(s-2)+2$ is for the sole reason that the numerator and squared expression in the Laplace transform of \(e^{-at}\sin t\) are both identical (i.e. \(s+a\) in numerator, \((s+a)^2\) in denominator). we're trying to make our expression look like Laplace transform templates that we recognize in order to invert the transform by eye

OpenStudy (anonymous):

awesome explanation!!! wow i want to give you a medal but I already gave it away! I will find you later and give you one on another thread! :o) that is such an awesome explanation! really...wow!

OpenStudy (anonymous):

if you instead split \(s=(s-2)+2\) you would still not be able to recognize any specific Laplace transform after breaking into two rational expressions, and inverting the Laplace transform in this manner is ultimately in recognizing the forms of common Laplace transforms

OpenStudy (anonymous):

I totally get it! :o)

OpenStudy (anonymous):

hey @ParthKohli ...give this gentleman a medal will you? :o)

OpenStudy (anonymous):

that being said, it's not a matter of *having* to do it in this way--there are ways of using Laplace properties to tackle this in a slightly different manner, which is generally the case. for example, as you noted: \[\mathcal{L}\left\{e^{-at}f(t)\right\}=F(s+a)\] where \(F(s)=\mathcal{L}\{f(t)\}\) in our case, we notice that we have a pesky \(s+2\) in our denominator, suggesting that we could apply the above Laplace transform property to rewrite our expression as follows: \[\mathcal{L}^{-1}\left\{\frac{(s+2)-2}{(s+2)^2+1}\right\}=e^{-2t}\cdot\mathcal{L}^{-1}\left\{\frac{s-2}{s^2+1}\right\}\]i.e. drawing out \(e^{-2t}\) in order to replace \(s+2\mapsto s\). now, you can split into two rational expressions naturally with the same result of \(e^{-2t}(\sin t-2\cos t)\)

OpenStudy (anonymous):

wow...do you know how to do L^-1{ 2s+5/s^2+6s+34} ?

OpenStudy (anonymous):

2S-5/(S+3)^2 +25 SO FAR

Parth (parthkohli):

May I try? :P\[\dfrac{2s+5}{(s+3)^2 + 25}=\dfrac{2(s+3) - 1}{(s+3)^2 + 25}\]Not sure if that works, but I just did what I saw you do in the previous question.

OpenStudy (anonymous):

\[L^-1(2S-5+3-3)/(S+3)^2+25\]

OpenStudy (anonymous):

how did you factor the 2?

Parth (parthkohli):

\[2s + 5 = 2s + 6 - 1 = 2(s+3) - 1\]

OpenStudy (anonymous):

did you do something other than +-3 ?

OpenStudy (anonymous):

nevermind...i see it...tricky

OpenStudy (anonymous):

that should work!...lemme finish :o)

OpenStudy (anonymous):

just to make a connection here, @ParthKohli is 'regrouping' in a way much like mental division; if you use polynomial division to evaluate \((2s+5)\div (s+3)\) you will find: $$\frac{2s+5}{s+3}=2\text{ with a remainder of }{-1}=2-\frac1{s+3}$$multiplying both sides by \(s+3\) gives \(2s+5=2(s+3)-1\)

OpenStudy (anonymous):

i get as far as e^-3tsin5t-e^-3t(something)

OpenStudy (anonymous):

so you're stuck at $$\mathcal{L}^{-1}\left\{-\frac{1}{(s+3)^2+25}\right\}$$?

OpenStudy (anonymous):

dont know

OpenStudy (anonymous):

well, recognize that we know: $$\mathcal{L}\{e^{-at}\cos(bt)\}=\frac{b}{(s+a)^2+b^2}$$; in this case we can see \(a=3\) and \(b^2=25\) suggesting we want \(b=5\). but our numerator looks like \(-1\) while we want it to look like \(5\)

OpenStudy (anonymous):

so \[-\frac1{(s+3)^2+5^2}\cdot\frac55=-\frac15\cdot\frac5{(s+3)^2+5^2}\] and then \[\mathcal{L}^{-1}\left\{-\frac15\cdot\frac5{(s+3)^2+5^2}\right\}=-\frac15\cdot\mathcal{L}^{-1}\left\{\frac5{(s+3)^2+5^2}\right\}=-\frac15e^{-3t}\cos(5t)\]

OpenStudy (anonymous):

actually I don't have that memorized...the teacher only made us remember 7 basic ones

OpenStudy (anonymous):

omg! that is seriously tricky algebra!

OpenStudy (anonymous):

well, do you have the following 'memorized'? $$\mathcal{L}\left\{\cos(bt)\right\}=\frac{b}{s^2+b^2}$$ what about this? $$\mathcal{L}\left\{e^{-at}f(t)\right\}=F(s+a)\quad\text{where }F(s)=\mathcal{L}\{f(t)\}$$

OpenStudy (anonymous):

so should I just keep in mind to make k^2 on the bottom and then think about fancy forms of 1?

OpenStudy (anonymous):

1st one yes...second one no, and my book has the signs reversed also

OpenStudy (anonymous):

in this case, we've just applied both at the same time :-) you can also apply the exponential property first as I demonstrated earlier. and yes, you should recognize the \(b^2\) in the denominator suggests our numerator needs to look like \(b\) and we can accomplish that using a 'fancy one', right

OpenStudy (anonymous):

the second identity can also be written by substituting \(a\mapsto-a\) which gives $$\mathcal{L}\left\{e^{at}f(t)\right\}=F(s-a)\quad\text{where }F(s)=\mathcal{L}\{f(t)\}$$

OpenStudy (anonymous):

so then how do i do 5/(s+3)^2 + 5^2 part then?

OpenStudy (anonymous):

to be completely honest with you, I'm rather aversive to memorizing properties and instead I typically derive both of those results on the fly from the integrals :)

OpenStudy (anonymous):

for that part, we'll notice we can map \(s+3\mapsto s\) if we pull out an \(e^{-3t}\) using the exponential property: $$\mathcal{L}\left\{\frac5{(s+3)^2+5^2}\right\}=e^{-3t}\cdot\mathcal{L}^{-1}\left\{\frac{5}{s^2+5^2}\right\}=e^{-3t}\cos(5t)$$

OpenStudy (anonymous):